OpenStudy (anonymous):

Write the standard form of the equation of the line described. #1) through: (-5,3), perpendicular to y=5/7x+4.

OpenStudy (anonymous):

do you know what the slope of \[\huge y=\color{red}{\frac{5}{7}}x+4\] is ?

OpenStudy (anonymous):

5/7

OpenStudy (anonymous):

ok good so the perpendicular line will have slope ?

OpenStudy (anonymous):

5/7

OpenStudy (anonymous):

no, that would be the slope of any parallel line you are asked for a perpendicular line

OpenStudy (anonymous):

what would it be?

OpenStudy (anonymous):

in english (math english) it is the "negative reciprocal" i.e. flip it and change the sign

OpenStudy (anonymous):

-7/5

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now use the point slope formula to find the equation, since you have a point and the slope

OpenStudy (anonymous):

y=-7/5x-5, is this the correct answer?

OpenStudy (anonymous):

i doubt it, since it has all those fractions

OpenStudy (anonymous):

\[y-3=-\frac{7}{5}(x+5)\] is a start

OpenStudy (anonymous):

multiply out and get \[y-3=-\frac{7}{5}x-7\] oh look the fractions are gone, but now you add \(3\) to both sides

OpenStudy (anonymous):

y=-7/5x-4

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

My textbook said the answer would be 7x+5y=-20

OpenStudy (anonymous):

that is because your text had you answer in standard from rather than point slope form

OpenStudy (triciaal):

the question asked for standard form